OpenStudy (anonymous):

hey can someone answer this super hard math problem? A wagon train that is one mile long advances one mile at a constant rate. During the same time period, the wagon master rides his horse at a constant rate from the front of the wagon train to the rear, and then back to the front. How far did that wagon master ride?

OpenStudy (anonymous):

this ur answ34

OpenStudy (anonymous):

i love those type of hats what are they called.. i think fedoras

OpenStudy (anonymous):

thi sme tryong to figure out math but im not mad about it x

OpenStudy (anonymous):

HAHA

OpenStudy (anonymous):

ya um idk

OpenStudy (anonymous):

can you please stop @ivanapen8

OpenStudy (anonymous):

I am trying to get an answer this is due tomorrow

OpenStudy (anonymous):

ok sure thing @daniellecrook

OpenStudy (anonymous):

got em

OpenStudy (anonymous):

tht nice

OpenStudy (anonymous):

ok thanks bean

OpenStudy (anonymous):

this look like the girl who helped me with my question i love her

OpenStudy (anonymous):

sweet life

OpenStudy (anonymous):

this tasted good

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20080701115439AAyLlXW its ok i dont need ur answers

OpenStudy (anonymous):

this was the soup

OpenStudy (anonymous):

thats our math teacher IM LAUGHING

OpenStudy (anonymous):

that looks really yummy mm make sure u dont cut yourself on that plate!

OpenStudy (anonymous):

HA

OpenStudy (anonymous):

matt just left the house

OpenStudy (anonymous):

these thinkgs arent good

OpenStudy (anonymous):

ok???????????

OpenStudy (anonymous):

tbh its been showing this for the last hour

OpenStudy (anonymous):

tht a no brainer

OpenStudy (anonymous):

just do it

OpenStudy (anonymous):

oldrin type

OpenStudy (anonymous):

send it

OpenStudy (anonymous):

wow r u kidding

OpenStudy (anonymous):

HAHAHHA

OpenStudy (anonymous):

oldrin is still typing the mathematician is hard at work

OpenStudy (anonymous):

Let's say the wagon masters absolute speed is \(r_m\), and so his effective (relative) speed towards the rear of the train is given by \(r_m+r_w\) (the train is approaching him). His relative speed towards the front of the train, however, is only \(r_m-r_w\) (the train is moving away from him). It should make sense that going towards the rear, our biker effectively travels \(1\) mile at a rate of \(r_m+r_w\) and hence it takes him \(\dfrac1{r_m+r_w}\). On the way back, however, he effectively travels \(1\) mile at a rate of \(r_m+r_w\) and it takes him \(\dfrac1{r_m-r_w}\). Consider, then, that total amount of time taken must be merely the time it took the wagon itself to travel that far i.e. \(\frac1{r_w}\) hence:$$\frac1{r_m+r_w}+\frac1{r_m-r_w}=\frac1{r_w}\\\frac{r_m-r_w+r_m+r_w}{r_m^2-r_w^2}=\frac1{r_w}\\2r_mr_w=r_m^2-r_w^2\\r_m^2-2r_mr_w-r_w^2=0\\r_m^2-2r_mr_w+r_w^2=2r_w^2\\(r_m-r_w)^2=2r_w^2\\r_m-r_w=r_w\sqrt2\\r_m=(1+\sqrt2)r_w$$Thus the master travels at the absolute rate of \((1+\sqrt2)r_w\) and so in the same time span (i.e. \(1/r_w\)) travels \((1+\sqrt2)r_w\times1/r_w=1+\sqrt2\) miles.

OpenStudy (anonymous):

typo: he effectively travels at a rate of \(r_m\color{red}-r_w\) on the way back to the head of the wagon train

OpenStudy (anonymous):

ty so much